OpenStudy (anonymous):

Let f(x) = x + 8 and g(x) = x2 - 6x - 7. Find f(g(2)). -7 -3 10 33

OpenStudy (bibby):

Start by plugging in 2 to g(x) 2^2 - (6*2) - 7

OpenStudy (anonymous):

okay now what

OpenStudy (bibby):

What'd you get? Plug what you get into f(x)

OpenStudy (anonymous):

f(g(x)) = f(x^2-6x-7) = x^2-6x -7 + 8 f(g(x)) =x^2 -6x + 1 So, f(g(2)) = (2)^2-6(2) +1 = 4-12 + 1 =5-12 = -7